OpenStudy (anonymous):

What are the amplitude, period, and phase shift of the given function? f(t) = 1/2 cos(3t - 2pi)

OpenStudy (tkhunny):

You MUST get the Amplitude. \(f(t) = a\cos(b(x-\alpha))\) a, b, and \(\alpha\) are important and they answer all your questions.

OpenStudy (anonymous):

so whats the answer?

OpenStudy (anonymous):

because i dont know how to do this

OpenStudy (tkhunny):

You had better study your course materials. We cannot teach the whole class, here. If you truly have no idea, you are likely in the wrong class and it's not worth your time to play with it. Are you int he right class? Have you been attending? Are you reading the materials?

OpenStudy (anonymous):

K